# Quantitative characterization of finite simple groups: a complement

### v3

@@ 2. The spectrum characterization of finite simple groups
Let $L$ be an unrecognizable simple group. Then $L$ is isomorphic to one of the following groups: $A_{6}$, $A_{10}$, $J_{2}$, ${}^{3}D_{4}(2)$, $L_{3}(3)$, $U_{3}(3)$, $U_{3}(5)$, $U_{3}(q)$ where $q\geqslant 7$ and $q$ is special Mersenne, $U_{5}(2)$, $S_{4}(3)$, $S_{4}(2^{m})$, $S_{4}(3^{m})$ with $m$ even, $S_{4}(q)$ where $q=p^{m}$ and $p\geqslant 5$, $S_{8}(q)$ for $q=p^{m}$ and $p\neq 2,7$, $O_{9}(q)$ where $q=p^{m}$.
@@
The spectrum of a finite group is a set of positive integers containing $1$. If the integer set $\pi_{e}(G)$ is partitioned into $1$, the set of prime number $\pi_{e}^{\prime}(G)$ ($\pi_{e}^{\prime}(G)=\pi(G)$) and the set of composite numbers $\pi^{\prime\prime}_{e}(G)$, then the following result holds:
@@
In the cases (II)–(VI) of the aforementioned theorem, is the number of simple groups satisfying the respective conditions finite or infinite?
@@
Notice that the order of an element in a group is the order of cyclic subgroup. Yanquan Feng [58] and Shi [188] provided finite groups with orders of abelian subgroups and orders of proper subgroups are consecutive integers, respectively. Furthermore, since the order of an element in a group is a conjugation invariant. Guohua Qian [167] investigated the finite groups with consecutive nonlinear character degrees.
@@ 3. Characterizing finite simple groups using the order of group and the set of element orders
To finally prove Conjecture 1.2, the nonabelian composition factors for these symplectic groups and orthogonal groups are given directly from the spectrum of the classical groups $B_{n}(q)$, $C_{n}(q)$, and $D_{n}(q)$ (where $n$ is even) in paper [211]. Finally, Conjecture 1.2 was proved in [212] by using the three theorems of [211]. Consequently, Conjecture 1.2, which characterizes finite simple groups by using “two orders”, has been proved and established as a theorem.
@@
Qian[168] and Qian, et al.[171] defined codegree of the character and gave a connection between the element order and the degree of character. Next, we consider the characterization of simple groups by using “the order of Hall subgroup of a group” and the “set of element orders”. Firstly, the “order ($|G|_{2}$) of Sylow $2$-subgroups” of a group $G$ and the “set of element orders” were used to characterize alternating and sporadic simple groups. We have the following theorems:
@@
The authors [140, 165, 180] earlier studied the characterization of alternating groups and symmetric groups using the set of element orders. The subsequent studies in [64, 117, 231] the authors further investigated the characterization of alternating groups by using the set of element orders. It was pointed out in [204] that $A_{6}$ and $A_{10}$ are unrecognizable.
@@
In 2007, Mazurov posed the following conjecture at International Algebraic Conference held in St. Petersburg: Let $G$ be a finite group, and let $L$ be an alternating group of sufficiently large degree or a simple group of Lie type of sufficiently large Lie rank. If $G$ is isospectral to $L$, then $G$ is an almost simple group with socle isomorphic to $L$. In fact, if $L$ is a finite simple group and $G$ is isospectral to $L$, then either $G$ is solvable, in which case $L$ is one of the following cases: $L_{3}(3)$, $U_{3}(3)$, and $S_{4}(3)$; or $G$ is an almost simple group with socle isomorphic to $L$ (See [64, Theorem 2]). The conjecture proposed by Mazurov was finally proved to be valid in [76]. By using these results and comparing $|G|_{\pi}$ and $|M|_{\pi}$, Question 3.18 can be answered.
@@ 4.1. The origin and early work of the Thompson Conjecture
Guiyun Chen [37] proved that if $M$ is a sporadic simpe group then Thompson conjecture is correct. Subsequently, he finished his PhD thesis titled “On Thompson Conjecture”. The author [38] based on the foundation of prime graph components in [219], a definition of the order components was provided as follows:
@@
Thompson Conjecture was proved to be ture for all sporadic simple groups (See [37, 38]). Its validity was also proved for finite group having at least $3$ prime graph components (See [39, 40]). For the case of $2$ prime graph components, we can use the same method to prove it. For instance, it was proved that Thompson Conjecture is correct for the group ${}^{3}D_{4}(q)$ (See [41]).
@@ 4.2. The case of groups with connected prime graph
In 2009, Vasil’ev proved that Thompson Conjecture is valid for the simple groups $A_{10}$ and $L_{4}(4)$ with connected prime graph. Subsequently, N. Ahanjideh proved that Thompson Conjecture is also correct for simple groups of Lie type $A_{n}(q)$ (See [5]), $B_{n}(q)$ (See [8]), $C_{n}(q)$ (See [6]), $D_{n}(q)$ where $n\neq 4,8$ (See [7]), ${}^{2}A_{n}(q)$ (See [9]) and ${}^{2}D_{n}(q)$ (See [10]) in a series of articles. For other exceptional simple groups of Lie type, Thompson Conjecture was established the validity in the literature [70, 94, 108, 109, 224]. The authors [13, 63, 65, 66, 67, 68, 126, 222] proved that Thompson Conjecture is true for alternating simple groups. Finally, I. B. Gorshkov [69] proved the validity of Thompson Conjecture for the remaining cases $D_{4}(q)$ and $D_{8}(q)$. Therefore, Thompson Conjecture was completely proved and was became a theorem.
@@ 4.4. Order equation
Let $M_{G}(n)$ be the set of elements of $G$ with order $n$, the elements of $G$ can be partitioned into the elements of same order, that is, $G=\bigcup M_{G}(n)$ where $n\in\pi_{e}(G)$. Since every element of order $n$ must be contained in some cyclic subgroup of order $n$, we have $|M_{G}(n)|=V_{n}(G)\phi(n)$, where $V_{n}(G)$ is the number of cyclic subgroups of order $n$, $\phi(n)$ is Euler totient function. Hence we can obtain the following equation, $|G|=\bigcup V_{n}(G)\phi(n)$, $n\in\pi_{e}(G)$. Here we denote the “order equatin” of $G$ as $Ord(G)$. That is, if two order equations of finite groups $G$ and $H$ are same, then $\pi_{e}(G)=\pi_{e}(H)$ and $V_{d}(G)=V_{d}(F)$ for any $d\in\pi_{e}(G)$. It is easy to see that same order equation is equivalent to the same order type of Thompson Problem.
Let $G$ and $H$ are two groups of same order type, it is not hard to prove that if $H$ is nilpotent then $G$ is also nilpotent and if $H$ is supersolvable then $G$ is solvable (See [221]). Of course, if $H$ is a finite nonabelian simple group, then $G\cong H$ is nonsolvable because Conjecture 1.2 has been proved. It was proved that if $H$ is solvable and the prime graph of $H$ is disconnected, then $G$ is solvable (See [173, 174]). The following results were proved in lecture [173, 174]:
@@
A group $G$ is isomorphic to $A_{5}$ if and only if the same order type of $G$ is $\{1,15,20,24\}$.
@@
$G_{1}$ and $G_{2}$ are of the same order type if and only if $|G_{1}(d)|=|G_{2}(d)|$, $d=1,2,\cdots$
Definition 4.9.
@@
The authors [124] proved that if $G_{1}$ and $G_{2}$ are of the same Burnside ring type, then $G_{1}$ and $G_{2}$ are of the same order type. Thus $G_{1}$ and $G_{2}$ are of the same two orders type immediately. But the reverse side is not correct. We have the following examples:
@@
Let $G_{1}=L_{3}(4):2_{2}$ and $G_{2}=2^{4}:A_{7}$. Then $G_{1}$ and $G_{2}$ are the same order type, but $B(G_{1})\not\cong B(G_{2})$.
@@
Let $G_{1}=A_{5}\times A_{5}\times A_{5}$, $G_{1}$ is not solvable and $|G_{1}|=60^{3}$ and $\pi_{e}(G_{1})=\{1,2,3,5,6,10,15,30\}$ since $\pi_{e}(A_{5})=\{1,2,3,5\}$. Let $G_{2}=C_{30}\times C_{30}\times C_{30}\times C_{2}\times C_{2}\times C_{2}$. Then $|G_{2}|=60^{3}$ and $\pi_{e}(G_{2})=\{1,2,3,5,6,10,15,30\}$. $G_{1}$ and $G_{2}$ are the same two orders type, but $G_{2}$ is solvable. Moreover, $G_{1}$ and $G_{2}$ are not the same order type.
@@ 5.1. Width of order and width of spectrum in finite groups
In August 2020, at the online conference “Ural Workshop on Group Theory and Combinatorics” held in Russia, the author gave the first presentation and introduced the aforementioned two definitions of width, and discussed the structure of finite groups with a small width, we refer to [196] for some results about the “width of order”. This literature points out that groups with $\omega_{0}(G)=1$, finite $p$-groups, although the orders of these groups have only one prime factor, their structures are quite complex. Recent Chinese literature can be found in [239, 240]. For the especially “simple” condition, $\pi_{e}(G)=\{1,3\}$, but its structure is quite complex, and currently, there is no a detailed classification.
@@
There are infinite simple $K_{4}$ groups.
@@
Are there infinite simple groups whose order being a square number?
@@ 6.1. Group and quantity
Let $G$ be a finite group and let $\pi_{e}(G)$ be the set of element orders in $G$. The set of numbers $T$ is called “set of intersection empty solvable” if $\pi_{e}(G)\cap T=\varnothing$ implies that $G$ is a solvable group.
@@
Let $G$ be a finite group and let $\pi_{e}(G)$ be the set of element orders in $G$. If $\pi_{e}(G)\cap T=\varnothing$ where $T=\{2\}$, $\{3,4\}$ or $\{3,5\}$, then $G$ is solvable. Furthermore, using the set of intersection empty solvable of $\pi_{e}(G)$ to judge $G$ is solvable or not, only the above three cases.
@@ 6.2. Group and graph are closely interconnected
Gruenberg and Kegel  [^5] K. W. Gruenberg, O. H. Kegel. Unpublished manuscript, 1975 first introduced the definition of prime graph $\Gamma(G)$ of group $G$ (See Definition 2.18) and its classification. Subsequently, the literature [93, 113, 114, 219] provided explicit descriptions for the connected components of the prime graphs of the finite simple groups. M. S. Lucido defined the diameter of the prime graph as follows (See [127]):
@@
If $G$ is a finite slvable group then $\operatorname{diam}(\Gamma(G))\leqslant 3$. Lucido [127] proved that for all finite groups $G$, we have $\operatorname{diam}(\Gamma(G))\leqslant 5$ and characterized almost simple groups with $\operatorname{diam}(\Gamma(G))=5$. Recall that a group $B$ is called an almost simple group if there exists a nonabelian simple group $A$ such that $A\leqslant B\leqslant\operatorname{Aut}(A)$.
@@ 6.4. Unrecognizable finite group by prime graph
Similar to the previous discussion about the function $h$ (See Definition 2.5). Maslova, V. V. Panshin and A. M. Staroletov [137] gave the following definition.
@@ 6.5. Graphs defined on groups
There are various graphs defined on a group, but the one most closely associated with quantitative characterization is the prime graph (GK graph). Cameron [31] introduced the following graphs: Their vertex sets are finite groups $G$, and their edges reflect the structure of the groups in a certain way, such that the automorphism group of $G$ replaces the automorphism of the graph. These graphs include the commuting graph (two vertices $x$ and $y$ are joined if and only if $xy=yx$; this graph was first introduced in 1955, see [26]), the generating graph ($x$ and $y$ joined if and only if $\langle x,y\rangle=G$; the generating graph was first introduced in 1996, see reference [27, 125]), the power graph (vertices $x$ and $y$ are joined if and only if one of $x$ and $y$ is a power of the other; this graph was first studied in 2000, see [103, 156, 118]), the enhanced power graph (two vertices $x$ and $y$ are joined if and only if $\langle x,y\rangle$ is not cyclic; this graph was first studied in 2007, see [2, 3]) and the deep commuting graph (two vertices, $x$ and $y$, are joined if and only if their preimages, $\overline{x}$ and $\overline{y}$, commute in every central extension of $G$. In other words, for every group $H$ with a central subgroup $Z$ such that $H/Z\cong G$, and their preimages $\overline{x}$ and $\overline{y}$ commute in $H$, see [32]).
This paper mainly discusses the quantitative characterization of finite simple groups, which is an essential component of quantitative properties of groups and a significant topic of group theory research. Studying finite groups using the set of element orders and characterizing finite simple groups by “two orders” were first proposed and investigated by the author, which have a significant influence on the field of group theory at home and abroad. Since the 1980s, the author has successively worked at Southwest Normal University (including a concurrent doctoral supervisor position at Sichuan University), Soochow University, and Chongqing University of Arts and Sciences. Additionally, the author has received funding for 11 projects from the National Natural Science Foundation of China, enabling continuous progress in this work. Our intention in writing this article is not only to systematically summarize our work, distill new viewpoints, ideas, and methods to solve related problems, but also to further delve into this work and find more significant applications.

### v5

@@ 2. The spectrum characterization of finite simple groups
For the case of $h(\pi_{e}(G))=2$, Buturlakin (See [248] ) proved that if $G$ and $H$ are finite simple groups such that $\pi_{e}(H)=\pi_{e}(G)$ and $H\ncong G$, then $\{G,H\}=\{S_{6}(2),O^{+}_{8}(2)\}$ or $\{G,H\}=\{O_{7}(3),O^{+}_{8}(3)\}$. Moreover, the author of [248] proved that there are no the case of $h(\pi_{e}(G))=3$ for three different finite simple groups $G$.
@@
Let $L$ be one of the following groups: $A_{6}$, $A_{10}$, $J_{2}$, ${}^{3}D_{4}(2)$, $L_{3}(3)$, $U_{3}(3)$, $U_{3}(5)$, $U_{3}(q)$ where $q\geqslant 7$ and $q$ is special Mersenne, $U_{5}(2)$, $S_{4}(3)$, $S_{4}(2^{m})$, $S_{4}(3^{m})$ with $m$ even, $S_{4}(q)$ where $q=p^{m}$ and $p\geqslant 5$, $S_{8}(q)$ for $q=p^{m}$ and $p\neq 2,7$, $O_{9}(q)$ where $q=p^{m}$ Then $L$ is an unrecognizable simple group.
@@
The spectrum of a finite group is a set of positive integers containing $1$. If the integer set $\pi_{e}(G)$ is partitioned into $1$, the set of prime numbers $\pi_{e}^{\prime}(G)$ and the set of composite numbers $\pi^{\prime\prime}_{e}(G)$, then the following result holds:
@@
In the cases (II)-(VI) of the aforementioned theorem, is the number of simple groups satisfying the respective conditions finite or infinite?
@@
Notice that the order of an element in a group is the order of cyclic subgroup. Yanquan Feng [58] and the author [188] provided finite groups with orders of abelian subgroups and orders of proper subgroups are consecutive integers, respectively. Furthermore, since the order of an element in a group is a conjugation invariant. Guohua Qian [167] investigated the finite groups with consecutive nonlinear character degrees.
@@ 3. Characterizing finite simple groups using the order of group and the set of element orders
To finally prove Conjecture 1.2, restrictions on nonabelian composition factors of finite groups have the same spectrum as $B_{n}(q)$, $C_{n}(q)$, and $D_{n}(q)$ (where $n$ is even) in [211]. Finally, Conjecture 1.2 was proved in [212] by using the three theorems of [211]. Consequently, Conjecture 1.2, which characterizes finite simple groups by using “two orders”, has been proved and established as a theorem.
@@
Qian[168] and Qian, et al.[171] defined codegree of the character and gave a connection between the element order and the degree of character.
Next, we consider the characterization of simple groups by using “the order of Hall subgroup of a group” and the “set of element orders”. Firstly, the “order ($|G|_{2}$) of Sylow $2$-subgroups” of a group $G$ and the “set of element orders” were used to characterize alternating and sporadic simple groups. We have the following theorems:
@@
The authors of [140, 165, 180] earlier studied the characterization of alternating groups and symmetric groups using the set of element orders. The subsequent studies in [64, 117, 231] the authors further investigated the characterization of alternating groups by using the set of element orders. It was pointed out in [204] that $A_{6}$ and $A_{10}$ are unrecognizable.
@@
In 2007, Mazurov posed the following conjecture at International Algebraic Conference held in St. Petersburg: Let $G$ be a finite group, and let $L$ be an alternating group of sufficiently large degree or a simple group of Lie type of sufficiently large Lie rank. If $G$ is isospectral to $L$, then $G$ is an almost simple group with socle isomorphic to $L$. In fact, if $L$ is a finite simple group and $G$ is isospectral to $L$, then either $G$ is solvable, in which case $L$ is one of the following cases: $L_{3}(3)$, $U_{3}(3)$, and $S_{4}(3)$; or $G$ has exactly one nonabelian composition factor (See [246], [214, Corollary 7.3] and [64, Theorem 2]). The conjecture proposed by Mazurov was finally proved to be valid in [76]. By using these results and comparing $|G|_{\pi}$ and $|M|_{\pi}$, Problem 3.18 can be answered.
@@ 4.1. The origin and early work of the Thompson Conjecture
Guiyun Chen [37] proved that if $M$ is a sporadic simpe group then Thompson conjecture is correct. Subsequently, he finished his PhD thesis titled “On Thompson Conjecture”. The author (See [38] ) based on the foundation of prime graph components in [219], a definition of the order components was provided as follows:
@@
Thompson Conjecture was proved to be true for all sporadic simple groups (See [37, 38]). Its validity was also proved for finite group having at least three prime graph components (See [39, 40]). For the case of two prime graph components, we can use the same method to prove it. For instance, it was proved that Thompson Conjecture is correct for the group ${}^{3}D_{4}(q)$ (See [41]).
@@ 4.2. The case of groups with connected prime graph
In 2009, Vasil’ev proved that Thompson Conjecture is valid for the simple groups $A_{10}$ and $L_{4}(4)$ with connected prime graph(See [210]). Subsequently, N. Ahanjideh proved that Thompson Conjecture is also correct for simple groups of Lie type $A_{n}(q)$ (See [5]), $B_{n}(q)$ (See [8]), $C_{n}(q)$ (See [6]), $D_{n}(q)$ where $n\neq 4,8$ (See [7]), ${}^{2}A_{n}(q)$ (See [9]) and ${}^{2}D_{n}(q)$ (See [10]) in a series of articles. For other exceptional simple groups of Lie type, Thompson Conjecture was established the validity in the literature [70, 94, 108, 109, 224]. The authors [13, 63, 65, 66, 67, 68, 126, 222] proved that Thompson Conjecture is true for alternating simple groups. Finally, I. B. Gorshkov [69] proved the validity of Thompson Conjecture for the remaining cases $D_{4}(q)$ and $D_{8}(q)$. Therefore, Thompson Conjecture was completely proved and was became a theorem.
@@ 4.4. Order equation
Let $M_{G}(n)$ be the set of elements of $G$ with order $n$, the elements of $G$ can be partitioned into the elements of same order, that is, $G=\bigcup M_{G}(n)$ where $n\in\pi_{e}(G)$. Since every element of order $n$ must be contained in some cyclic subgroup of order $n$, we have $|M_{G}(n)|=V_{n}(G)\phi(n)$, where $V_{n}(G)$ is the number of cyclic subgroups of order $n$, $\phi(n)$ is Euler totient function. Hence we can obtain the following equation, $|G|=\Sigma V_{n}(G)\phi(n)$, $n\in\pi_{e}(G)$. Here we denote the “order equation” of $G$ as $Ord(G)$. That is, if two order equations of finite groups $G$ and $H$ are same, then $\pi_{e}(G)=\pi_{e}(H)$ and $V_{d}(G)=V_{d}(F)$ for any $d\in\pi_{e}(G)$. It is easy to see that same order equation is equivalent to the same order type of Thompson Problem.
Let $G$ and $H$ are two groups of same order type, it is not hard to prove that if $H$ is nilpotent then $G$ is also nilpotent and if $H$ is supersolvable then $G$ is solvable (See [221]). Of course, if $H$ is a finite nonabelian simple group, then $G\cong H$ is nonsolvable because Conjecture 1.2 has been proved. It was proved that if $H$ is solvable and the prime graph of $H$ is disconnected, then $G$ is solvable (See [173, 174]). The following results were proved in papers [173, 174]:
@@
A group $G$ is isomorphic to $A_{5}$ if and only if $G$ has the same order type as $A_{5}$ ( the order type of $A_{5}$ is {1, 15, 20, 24 } ).
@@
The authors [124] proved that if $G_{1}$ and $G_{2}$ are of the same Burnside ring type, then $G_{1}$ and $G_{2}$ are of the same order type (see Def. 1.3). Thus $G_{1}$ and $G_{2}$ are of the same two orders type immediately. But the reverse side is not correct. We have the following examples:
@@
Let $G_{1}=L_{3}(4):2_{2}$ and $G_{2}=2^{4}:A_{7}$. Then $G_{1}$ and $G_{2}$ are of the same order type, but $B(G_{1})\not\cong B(G_{2})$.
@@
Let $G_{1}=A_{5}\times A_{5}\times A_{5}$, $G_{1}$ is not solvable and $|G_{1}|=60^{3}$ and $\pi_{e}(G_{1})=\{1,2,3,5,6,10,15,30\}$ since $\pi_{e}(A_{5})=\{1,2,3,5\}$. Let $G_{2}=C_{30}\times C_{30}\times C_{30}\times C_{2}\times C_{2}\times C_{2}$. Then $|G_{2}|=60^{3}$ and $\pi_{e}(G_{2})=\{1,2,3,5,6,10,15,30\}$. $G_{1}$ and $G_{2}$ are of the same two orders type, but $G_{2}$ is solvable. Moreover, $G_{1}$ and $G_{2}$ are of not the same order type.
@@
Recently, the author of [247] construct two finite groups of size $2^{365}\textperiodcentered 3^{105}\textperiodcentered 7^{104}$: a solvable group $G$ and a non-solvable group $H$, such that for every integer $n$ the groups have the same number of elements of order $n$. It gives a negative answer to Thompson problem. As a much smaller counterexample to Thompson problem, reference [249] constructed two same order type groups of order $2^{13}\textperiodcentered 3^{4}\textperiodcentered 7^{3}$, one solvable and the other non-solvable.
@@ 5.1. Width of order and width of spectrum in finite groups
In August 2020, at the online conference “Ural Workshop on Group Theory and Combinatorics” held in Russia, the author gave the first presentation and introduced the aforementioned two definitions of width, and discussed the structure of finite groups with a small width, we refer to [196] for some results about the “width of order”. This paper points out that groups with $\omega_{0}(G)=1$, finite $p$-groups, although the orders of these groups have only one prime factor, their structures are quite complex. Recent Chinese literature can be found in [239, 240]. For the especially “simple” condition, $\pi_{e}(G)=\{1,3\}$, but its structure is quite complex, and currently, there is no a detailed classification.
@@
There are infinitely simple $K_{4}$ groups.
@@
Are there infinitely simple groups whose order being a square number?
@@ 5.2. Finite groups with a small spectrum width
About Theorem 5.7 the authors of [33] have summerized in their paper all known results about non-solvable EPPO-groups in the form ”if and only if” (See [33, Theorem 1.7]).
@@ 6.1. Group and quantity
Let $G$ be a finite group and let $\pi_{e}(G)$ be the set of element orders in $G$. The set $T$ of numbers is called the set of ”intersection empty solvability” if $\pi_{e}(G)\cap T=\varnothing$ implies that $G$ is a solvable group.
@@
Let $G$ be a finite group and let $\pi_{e}(G)$ be the set of element orders in $G$. If $\pi_{e}(G)\cap T=\varnothing$ where $T=\{2\}$, $\{3,4\}$ or $\{3,5\}$, then $G$ is solvable. Furthermore, using the set T of ”intersection empty solvability” to judge $G$ is solvable or not, only the above three cases ( $T=\{2\}$, $\{3,4\}$ or $\{3,5\}$).
@@ 6.2. Group and graph are closely interconnected
Gruenberg and Kegel  [^5] K. W. Gruenberg, O. H. Kegel. Unpublished manuscript, 1975 first introduced the definition of prime graph $\Gamma(G)$ of group $G$ (See Definition 2.18) and its classification. Subsequently, the literature [93, 113, 114, 219] provided explicit descriptions for the connected components of the prime graphs of the finite simple groups. M. S. Lucido studied the diameter of the prime graph as follows (See [127]):
@@
If $G$ is a finite solvable group then $\operatorname{diam}(\Gamma(G))\leqslant 3$. Lucido [127] proved that for all finite groups $G$, we have $\operatorname{diam}(\Gamma(G))\leqslant 5$ and characterized almost simple groups with $\operatorname{diam}(\Gamma(G))=5$. Recall that a group $B$ is called an almost simple group if there exists a nonabelian simple group $A$ such that $A\leqslant B\leqslant\operatorname{Aut}(A)$.
@@ 6.4. Unrecognizable finite group by prime graph
Similar to the previous discussion about the function $h$ (See Definition 2.5). In [33] the authors gave the following definition (see [33], page 188).
@@ 6.5. Graphs defined on groups
There are various graphs defined on a group, but the one most closely associated with quantitative characterization is the prime graph (GK graph). Cameron [31] introduced the following graphs: Their vertex sets are finite groups $G$, and their edges reflect the structure of the groups in a certain way, such that the automorphism group of $G$ related to the automorphism of the graph. These graphs include the commuting graph (two vertices $x$ and $y$ are joined if and only if $xy=yx$; this graph was first introduced in 1955, see [26]), the generating graph ($x$ and $y$ joined if and only if $\langle x,y\rangle=G$; the generating graph was first introduced in 1996, see reference [27, 125]), the power graph (vertices $x$ and $y$ are joined if and only if one of $x$ and $y$ is a power of the other; this graph was first studied in 2000, see [103, 156, 118]), the enhanced power graph (two vertices $x$ and $y$ are joined if and only if $\langle x,y\rangle$ is not cyclic; this graph was first studied in 2007, see [2, 3]) and the deep commuting graph (two vertices, $x$ and $y$, are joined if and only if their preimages, $\overline{x}$ and $\overline{y}$, commute in every central extension of $G$. In other words, for every group $H$ with a central subgroup $Z$ such that $H/Z\cong G$, and their preimages $\overline{x}$ and $\overline{y}$ commute in $H$, see [32]).
This paper mainly discusses the quantitative characterization of finite simple groups, which is an essential component of quantitative properties of groups and a significant topic of group theory research. Studying finite groups using the set of element orders and characterizing finite simple groups by “two orders” were first proposed and investigated by the author, which have a significant influence on the field of group theory at home and abroad. Since the 1980s, the author has successively worked at Southwest Normal University (including a concurrent doctoral supervisor position at Sichuan University), Suzhou University, and Chongqing University of Arts and Sciences. Additionally, the author has received funding for 11 projects from the National Natural Science Foundation of China, enabling continuous progress in this work. Our intention in writing this article is not only to systematically summarize our work, distill new viewpoints, ideas, and methods to solve related problems, but also to further delve into this work and find more significant applications.